# Refined Bitcoin Security-Latency Under Network Delay

### v2

@@ Abstract
We study security-latency bounds for Nakamoto consensus, i.e., how secure a block is after it becomes $k$-deep in the chain. We improve the state-of-the-art bounds by analyzing the race between adversarial and honest chains in three different phases. We find the probability distribution of the growth of the adversarial chains under models similar to those in [1] when a target block becomes $k$-deep in the chain. We analyze certain properties of this race to model each phase with random walks that provide tighter bounds than the existing results. Combining all three phases provides novel upper and lower bounds for blockchains with small $\lambda\Delta$.
@@ 1 Introduction
Recently, several studies have given extensive bounds on the security-latency trade-off under network delay $\Delta$ [5, 1, 6]. They investigate how secure a block is at any point during the execution of the protocol under a network delay of $\Delta$. Specifically, [5] considers races between Poisson and renewal processes to give upper and lower bounds on how secure a block is after it is confirmed. [6] analyzes a dynamic programming algorithm to numerically bound the safety violation probability. While [5] considers confirmation rules that treat latency in terms of time units, [1] considers the case where the committing rule is defined in terms of how deep a block is in the current longest chain, and outperforms the results of [6] in small $\lambda\Delta$ regimes, such as Bitcoin where $\lambda\Delta\approx 1/60$.
@@
For Bitcoin, where a block is mined approximately every $10$ minutes and assuming network delay is at max $10$ seconds, under widely adopted $6$-block confirmation rule, safety violation probability is narrowed down to between $0.112\%$ and $0.173\%$ under $10\%$ adversarial presence. For comparison, the previous best-known results [1] were between $0.106\%$ and $0.353\%$.
@@ 2.3 Parameters and Definitions
(Rigged model) Some of the selected blocks mined by the honest nodes are converted into adversarial blocks, in addition to the blocks mined by the adversarial nodes themselves. Only the remaining blocks mined by the honest nodes are honest blocks [1].
@@
(Block publication) We say a block is published when it enters the view of every honest miner. As the adversary controls the network delay, in this context, a block published $\Delta$-delay after its mining time, $\tau_{m}$, means that the block enters the view of all honest nodes (except the miner) at $\tau_{m}+\Delta$.
@@
$\rho$ $\alpha\beta e^{-\lambda\Delta}$
$\rho^{\prime}$ $\alpha e^{-\lambda\Delta}(1+\lambda\Delta+\beta-\alpha e^{-\lambda\Delta})$
$\bar{\beta}^{2}$ $1-\bar{\alpha}^{2}-\rho^{\prime}$
@@ 3 Lower Bound
There are different achievable adversarial strategies. The strategy considered in [1] makes use of $\Delta=0$ which is a lower bound to non-zero delay since the adversary can always choose not to introduce any delays. Depending on the magnitude of $\lambda\Delta$, however, this bound may actually be quite off from reality. Here, we deploy an adversarial strategy known as private attack with $\Delta$ delay, a simple yet effective one to hamper the growth of the honest chain. Specifically, adversary delays every honest block by the maximum allowed $\Delta$, in the meanwhile, mining a private chain in order to double spend. In this scenario, the length of the honest chain will be equal to the number of published jumpers from the start of the protocol until the current time. Hence, under the $k$-block confirmation rule, we calculate a lower bound for the probability that the $(k-1)$st jumper after the target block is mined and at some point from then on $tx$ is excluded from the longest chain of an honest miner. Our analysis extends the one in [1] by incorporating the effects of $\Delta$ delay.
@@ 3.2 Confirmation Interval
At time $\tau$, target transaction $tx$ enters the transaction pool of miners and will be included in the next honest (jumper) block by assumption. We call the interval starting at $\tau$ and ending when the target block containing $tx$ becomes $k$-deep in the longest chain of all honest miners, the confirmation interval. Our goal is to find the number of adversarial blocks mined during this confirmation interval.
First, note that the lead found previously is a steady-state scenario and assumed to be independent of the Poisson arrivals starting at $\tau$, i.e., $\tau$ is large. Second, statistically, it does not matter if the confirmation interval ends when $tx$ becomes $k$-deep in any honest view or all honest views under the $\Delta$ delay strategy.
To see this second point, let $\tau_{c}$ denote the mining time of the $k$th jumper block mined starting from $\tau$. Then, the former interval (i.e., the interval when $tx$ becomes $k$-deep in any honest view) refers to $[\tau,\tau_{c}]$ whereas the latter interval (i.e., the interval when $tx$ becomes $k$-deep in all honest views) refers to $[\tau,\tau_{c}+\Delta]$. If the adversarial power can discard the target $tx$ in the former interval, it can do so in the latter interval as well, since the latter interval contains the former interval. If discarding happens in the latter interval, but not in the former interval, then, this necessarily implies that the adversary has mined some blocks during the $\Delta$ delay of the $k$th jumper block to catch the longest honest chain. However, if we were to use the former definition, we would have to consider the $\Delta$ delay interval of the $k$th jumper block separately right after the confirmation interval, since only the adversarial chain can grow during the $\Delta$ delay of the $k$th jumper block, ending up with the same result only to make the computations unnecessarily complex. Hence, we use the interval $[\tau,\tau_{c}+\Delta]$ as our confirmation interval.
@@
Let $C$ denote the number of adversarial blocks mined between the publication times of two consecutive jumper blocks. Next, Lemma 1 finds the PMF of $C$.
@@
Let $S$ denote the number of adversarial arrivals during the confirmation interval. Next, Corollary 1 finds the PMF of $S$ using the PMF of $C$ found in Lemma 1.
@@
Proof:  All jumpers including the first has the distribution of $C$ in (13). Further, $S$ is the sum of $k$ i.i.d. random variables each distributed as $C$. First, we find the probability generating function (PGF) of $C$, i.e., $G_{C}(z)=\mathbb{E}[z^{C}]$, as
@@ 3.3 Post-Confirmation Race
Proof:  We start by analyzing the case $n=1$. Let $B_{i}$ denote the position after the $i$th coin toss. Let $I$ denote the event that $\max\limits_{i\geq 1}B_{i}\geq m$. Note that $\max\limits_{i\geq 1}B_{i}$ has geometric distribution [9, eqn. (7.3.5)]. Further, given the event $I$, let $s$ denote the time of the first arrival on $m$ and $J_{s}$ denote the event that $\max\limits_{i>s}B_{i}<m$. Then, $\mathbb{P}(N=1|M=m)$ can be expressed as
@@
Proof:  Since by definition $T_{i}\leq T^{\prime}_{i}\leq T_{i}+1$, we have
@@
Let $\mathbb{P}(0\rightarrow 1)=p$ denote the probability of reaching state $1$ starting from the origin in a three-way coin toss model. Using the Chapman-Kolmogorov equations,
@@ 4 Upper Bound
To find an upper bound on the security-latency guarantee, one has to know the best adversarial strategy. It is proven in [7] that there is no deterministic adversarial strategy that outperforms other strategies for all possible arrivals under non-zero network delay; specifically, see [7, Lemma F. 1]. However, it is also proven in [1] that private attack is the best attack under the condition that all honest blocks are on different heights (we denote this condition in short as AHBODH). To make use of this fact, [1] uses a rigged model where some honest arrivals are converted into adversarial ones. The bound performs well for small $\lambda\Delta$ and $k$, however, as these parameters grow, the gap between the upper and lower bounds grows significantly. We improve this model by considering two arrivals at a time during pre-mining and post-confirmation intervals instead of one arrival at a time. We also use the Pascal distribution idea during the confirmation interval which gives an improvement.
@@ 4.1 Pre-Mining Gain
We will use this idea in the confirmation interval, however, we will calculate the lead with another method. For the sake of completeness, we also provide the steady state of $\bar{P}$ using Ramaswami’s formula in Appendix A. Here, we will consider two arrivals at a time during pre-mining which brings improvement to the lead considered in [1]. In the rigged model of [1], an arrival is i.i.d. Bernoulli, and coin toss lands honest w.p. $\alpha e^{-\lambda\Delta}$. Let $(H,A)$ denote the growth of honest and adversarial chains after two arrivals. If we consider two consecutive i.i.d. coin tosses separately and sum their results, then, 1. $\bar{E}_{1}^{\prime}$denoting $(2,0)$ happens w.p. $(\alpha e^{-\lambda\Delta})^{2}$; 2. $\bar{E}_{2}^{\prime}$denoting $(1,1)$ happens w.p. $2(\alpha e^{-\lambda\Delta})(1-\alpha e^{-\lambda\Delta})$; and 3. $\bar{E}_{3}^{\prime}$denoting $(0,2)$ happens w.p. $(1-\alpha e^{-\lambda\Delta})^{2}$. In our analysis $\bar{E}_{1}^{\prime}$ stays the same but we decrease the probability of $\bar{E}_{3}^{\prime}$ by considering two consecutive coin tosses together instead of considering them separately.
Assume that we start from time zero and consider groups of two arrivals in the rigged model. Note that the genesis block is the zeroth jumper that arrives at time zero. Here, $\bar{E}_{1}$, i.e., $(H,A)=(2,0)$ has the same probability as $\bar{E}_{1}^{\prime}$. For $\bar{E}_{2}$ to happen, i.e., $(H,A)=(1,1)$, there are $3$ cases:
  1. 1.
  2. 2.
  3. 3.
After the second arrival at $t_{2}$, by the memorylessness property of exponential arrivals, same arguments will hold for the next group of two arrivals and so forth. Hence, each group is identically distributed and independent from each other. Note that, $\mathbb{P}(\bar{E}_{2})=\rho^{\prime}\geq\mathbb{P}(\bar{E}_{2}^{\prime})$, hence the improvement. Putting these observations into a random walk starting from zero and moving with $(A,H)$, we obtain the following Markov chain,
$\!\!\!\bar{P^{\prime}}=\begin{bmatrix}1-\rho-\bar{\beta}^{2}&\rho&\bar{\beta}^{2}&0&0&0&0&\ldots\\ \bar{\alpha}^{2}&\rho^{\prime}&0&\bar{\beta}^{2}&0&0&0&\ldots\\ \bar{\alpha}^{2}&0&\rho^{\prime}&0&\bar{\beta}^{2}&0&0&\ldots\\ 0&\bar{\alpha}^{2}&0&\rho^{\prime}&0&\bar{\beta}^{2}&0&\ldots\\ 0&0&\bar{\alpha}^{2}&0&\rho^{\prime}&0&\bar{\beta}^{2}&\ldots\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\ddots\\ \end{bmatrix}$ (56)
where $\mathbb{P}(\bar{E}_{1})=\bar{\alpha}^{2}$, $\mathbb{P}(\bar{E}_{2})=\rho^{\prime}$ and $\mathbb{P}(\bar{E}_{3})=\bar{\beta}^{2}$. The resulting steady state distribution is $\bar{\pi}_{0}=\frac{\bar{\alpha}^{2}-\bar{\beta}^{2}}{\bar{\alpha}^{2}+\rho}$, $\bar{\pi}_{1}=\bar{\pi}_{0}\frac{\rho}{\bar{\alpha}^{2}}$ and
$\bar{\pi}_{i+2}=\bar{\pi}_{i}\left(\frac{\bar{\beta}}{\bar{\alpha}}\right)^{2}$ (57)
which is the distribution of the lead, $\bar{L^{\prime}}$. We denote the PMF of $\bar{L^{\prime}}$ with $P^{\prime}_{1}$.
Note that similar to a birth-death process, we cannot move left while at state zero. It is clear that the lead goes from zero to $2$ in $(1,3)$rd element of $\bar{P^{\prime}}$ with event $\bar{E}_{3}$. To understand why $(1,2)$nd element of $\bar{P^{\prime}}$ is $\rho$, first note that when the lead is zero, both cases 2 and 3 in $\bar{E}_{2}$ do not change the lead since the adversarial block arrives before the jumper block. Now, consider the mining events when the lead is zero and consider all subevents of case 1 in $\bar{E}_{2}$.
  1. 1.
    The next 2 arrivals could be honest and adversarial, respectively, with $t_{1}>\Delta$. In this case the lead increases from zero to $1$. This case has probability $\rho$.
  2. 2.
    The next 2 arrivals could be both honest with $t_{1}>\Delta$ and $t_{2}-t_{1}<\Delta$. In this case, we can assume that the first arrival is converted instead of the second one (which is also an upper bound for the adversary since it is not within physical capabilities of the adversary to convert a jumper). Then, the second arrival automatically becomes a jumper since $t_{2}>t_{1}>\Delta$ and AHBODH is still satisfied and the lead stays zero.

### v3

@@ Abstract
We study security-latency bounds for Nakamoto consensus, i.e., how secure a block is after it becomes $k$-deep in the chain. We improve the state-of-the-art bounds by analyzing the race between adversarial and honest chains in three different phases. We find the probability distribution of the growth of the adversarial chains under models similar to those in [Guo, Ren; AFT 2022] when a target block becomes $k$-deep in the chain. We analyze certain properties of this race to model each phase with random walks that provide tighter bounds than the existing results. Combining all three phases provides novel upper and lower bounds for blockchains with small $\lambda\Delta$.
@@ 1 Introduction
Recently, several studies have given extensive bounds on the security-latency trade-off under network delay $\Delta$ [4, 5, 6]. They investigate how secure a block is at any point during the execution of the protocol under a network delay of $\Delta$. Specifically, [4] considers races between Poisson and renewal processes to give upper and lower bounds on how secure a block is after it is confirmed. [6] analyzes a dynamic programming algorithm to numerically bound the safety violation probability. While [4] considers confirmation rules that treat latency in terms of time units, [5] considers the case where the committing rule is defined in terms of how deep a block is in the current longest chain, and outperforms the results of [6] in small $\lambda\Delta$ regimes, such as Bitcoin where $\lambda=1/600$ blocks/s and $\Delta\approx 10$ seconds according to the historical data.
@@
For Bitcoin, where a block is mined approximately every $10$ minutes and assuming network delay is at max $10$ seconds, under widely adopted $6$-block confirmation rule, safety violation probability is narrowed down to between $0.112\%$ and $0.177\%$ under $10\%$ adversarial presence. For comparison, the previous best-known results [5] were between $0.106\%$ and $0.353\%$.
@@ 2.3 Parameters and Definitions
(Rigged block) A block mined by an honest (adversarial) node is simply called an honest (adversarial) block. In this sense, a rigged block is an honest block that is treated as if it is an adversarial block [5]. The adversary can use the rigged block in any way it wants in order to increase the chance of double spend.
@@
(Block publication) We say a block becomes public when it enters the view of every honest miner. Any honest block mined at $\tau_{m}$, becomes public by $\tau_{m}+\Delta$, i.e., enters the view of all honest nodes by $\tau_{m}+\Delta$.
@@
$\rho$ $\alpha e^{-\lambda\Delta}(1+\lambda\Delta+\beta-\alpha e^{-\lambda\Delta})$
$\bar{\beta}^{2}$ $1-\bar{\alpha}^{2}-\rho$
$\bar{a}_{i}$ $\sum_{j\geq i}\bar{\alpha}_{j}+\beta*\mathbbm{1}_{i\leq 2}$
$\bar{b}_{i}$ $\sum_{j\geq i}\bar{\alpha}_{j}+\beta*\mathbbm{1}_{i\leq 1}$
$a_{i}$ $\sum_{j\geq i}\alpha_{j}+\beta*\mathbbm{1}_{i\leq 2}$
$b_{i}$ $\sum_{j\geq i}\alpha_{j}+\beta*\mathbbm{1}_{i\leq 1}$
@@ 3 Lower Bound
There are different achievable adversarial strategies. The strategy considered in [5] makes use of $\Delta=0$ which is a lower bound to non-zero delay since the adversary can always choose not to introduce any delays. Depending on the magnitude of $\lambda\Delta$, however, this bound may actually be quite off from reality. Here, we deploy an adversarial strategy known as private attack with $\Delta$ delay, a simple yet effective one to hamper the growth of the honest chain. Specifically, adversary delays every honest block by the maximum allowed $\Delta$, in the meanwhile, mining a private chain in order to double spend. As the adversary controls the network delay, in this context, a block published $\Delta$-delay after its mining time, $\tau_{m}$, means that the block becomes public exactly at $\tau_{m}+\Delta$.
In this scenario, the length of the honest chain will be equal to the number of published jumpers from the start of the protocol until the current time. Hence, under the $k$-block confirmation rule, we calculate a lower bound for the probability that the $(k-1)$st jumper after the target block is mined and at some point from then on $tx$ is excluded from the longest chain of an honest miner. Our analysis extends the one in [5] by incorporating the effects of $\Delta$ delay.
@@ 3.2 Confirmation Interval
At time $\tau$, which is assumed to be large and random, target transaction $tx$ enters the transaction pool of miners and will be included in the next honest block by assumption. However, $\tau$ may be in the $\Delta$ delay immediately following a jumper that arrived before $\tau$. In this case, since we assume that the adversary applies the $\Delta$-delay strategy, if an honest block containing $tx$ is mined during this $\Delta$-delay, it forks the last jumper of pre-mining phase and becomes void. Thus, $tx$ is effectively included in the honest chain after the end of the $\Delta$ delay immediately following the last jumper of the pre-mining phase. Hence, without loss of generality, we can assume that $tx$ enters the mempool at a time $\tau$, where the first honest block to be mined contains the $tx$ and is a jumper. Further, this observation implies that the steady-state distribution of $P^{\prime}$ is a lower bound for the adversarial lead. We call the interval starting at $\tau$ and ending when the target block containing $tx$ becomes $k$-deep in the longest chain of all honest miners, the confirmation interval. Our goal is to find the number of adversarial blocks mined during this confirmation interval.
First, note that the lead found previously is a steady-state scenario and assumed to be independent of the Poisson arrivals starting at $\tau$, i.e., $\tau$ is large. Second, statistically, it does not matter if the confirmation interval ends when $tx$ becomes $k$-deep in any honest view or all honest views under the $\Delta$ delay strategy. To see this second point, let $\tau_{c}$ denote the mining time of the $k$th jumper block mined starting from $\tau$. Then, the former interval (i.e., the interval when $tx$ becomes $k$-deep in any honest view) refers to $[\tau,\tau_{c}]$ whereas the latter interval (i.e., the interval when $tx$ becomes $k$-deep in all honest views) refers to $[\tau,\tau_{c}+\Delta]$. If the adversarial power can discard the target $tx$ in the former interval, it can do so in the latter interval as well, since the latter interval contains the former interval. If discarding happens in the latter interval, but not in the former interval, then, this necessarily implies that the adversary has mined some blocks during the $\Delta$ delay of the $k$th jumper block to catch the longest honest chain. However, if we were to use the former definition, we would have to consider the $\Delta$ delay interval of the $k$th jumper block separately right after the confirmation interval, since only the adversarial chain can grow during the $\Delta$ delay of the $k$th jumper block, ending up with the same result only to make the computations unnecessarily complex. Hence, we use the interval $[\tau,\tau_{c}+\Delta]$ as our confirmation interval.
@@
Let $C$ denote the number of adversarial blocks mined between the publication times of two consecutive jumper blocks. In Lemma 1 below, we find the PMF of $C$ by first considering the number of adversarial arrivals before the arrival of the honest block and add it to the number of adversarial arrivals during the $\Delta$ delay interval.
@@
Let $S$ denote the number of adversarial arrivals during the confirmation interval. Next, Corollary 1 finds the PMF of $S$ using the PMF of $C$ found in Lemma 1. To do so, we first express $C$ in terms of its probability generating function (PGF), note that PGF of $S$ is the $k$th power of the PGF of $C$, and finally transform the result back to the time domain.
@@
Proof:  All jumpers including the first has the distribution of $C$ in (3). Further, $S$ is the sum of $k$ i.i.d. random variables each distributed as $C$. First, we find the PGF of $C$, i.e., $G_{C}(z)=\mathbb{E}[z^{C}]$, as
@@ 3.3 Post-Confirmation Race
Proof:  We start by analyzing the case $n=1$, and observe that for $N=n+1$, the result is recursive by the Markov property. Let $B_{i}$ denote the position after the $i$th coin toss. Let $I$ denote the event that $\max\limits_{i\geq 1}B_{i}\geq m$. Note that $\max\limits_{i\geq 1}B_{i}$ has geometric distribution [9, eqn. (7.3.5)]. Further, given the event $I$, let $s$ denote the time of the first arrival on $m$ and $J_{s}$ denote the event that $\max\limits_{i>s}B_{i}<m$. Then, $\mathbb{P}(N=1|M=m)$ can be expressed as
@@
Proof:  We first find the distribution of $\max\limits_{i\geq 1}T_{i}$, which is geometric. Next, to consider the event of $E_{2}$ while $\max\limits_{i\geq 1}T_{i}=D-1$, we use the result of Corollary 2, which gives the distribution of the number of times $\max\limits_{i\geq 1}T_{i}=D-1$. Finally, we consider whether $E_{2}$ happened or not each time the walk hits $\max\limits_{i\geq 1}T_{i}=D-1$, and combine the results to complete the proof.
Since by definition $T_{i}\leq T^{\prime}_{i}\leq T_{i}+1$, we have
@@
Let $\mathbb{P}(0\rightarrow 1)=p$, where $0\rightarrow 1$ denotes the event of reaching state $1$ starting from the origin in a three-way coin toss model. Using the Chapman-Kolmogorov equations,
@@ 4 Upper Bound
To find an upper bound on the security-latency guarantee, one has to know the best adversarial strategy. It is proven in [7] that there is no deterministic adversarial strategy that outperforms other strategies for all possible arrivals under non-zero network delay; specifically, see [7, Lemma F. 1]. However, it is also proven in [5] that private attack is the best attack under the condition that all honest blocks are on different heights (we denote this condition in short as AHBODH). To make use of this fact, [5] uses a rigged model where some honest arrivals are converted into adversarial ones. The bound performs well for small $\lambda\Delta$ and $k$, however, as these parameters grow, the gap between the upper and lower bounds grows significantly. We refine this model by decreasing the number of rigged blocks without violating the AHBODH condition. We first modify $P$ matrix to incorporate the effect of the rigged blocks for the calculation of the lead. Next, we modify Lemma 1 and Corollary 1 to incorporate the effect of the rigged model during the confirmation interval. Finally, for the post-confirmation race, we consider two arrivals at a time which brings improvement to the model of [5].
@@ 4.1 Pre-Mining Gain
The steady state of matrix $\bar{P}$ can be found by Ramaswami’s formula [8], as $\bar{P}$ is a stochastic matrix of M/G/1 type.
Lemma 4
Steady state of $\bar{P}$ can be found recursively using
where $\pi_{0}=\frac{1-2\beta-\alpha\lambda\Delta}{\alpha}$.
Proof:  Note that, finding $\bar{a}_{i}$ and $\bar{b}_{i}$ is a straightforward application of Ramaswami’s formula and they are provided in Table 1. Here, we only prove $\pi_{0}=\frac{1-2\beta-\alpha\lambda\Delta}{\alpha}$, which is done by summing each hand side of the Ramaswami equation for $i\geq 1$ in order to express $\pi_{i}$ in terms of $\pi_{0}$.
Summing (35) for $i\geq 1$, we get
$(1-\bar{a}_{1})\sum_{i\geq 1}\pi_{i}$ $=\sum_{i\geq 1}\left(\pi_{0}\bar{b}_{i}+\sum_{j=1}^{i-1}\pi_{j}\bar{a}_{i+1-j}\right)$ (36)
$=\pi_{0}\sum_{i\geq 1}\bar{b}_{i}+\sum_{i\geq 2}\sum_{j=1}^{i-1}\pi_{j}\bar{a}_{i+1-j}$ (37)
$=-\pi_{0}\beta+\pi_{0}\sum_{i\geq 1}\bar{a}_{i}+\sum_{j\geq 1}\sum_{i>j}\pi_{j}\bar{a}_{i+1-j}$ (38)
$=-\pi_{0}\beta+\pi_{0}\sum_{i\geq 1}\bar{a}_{i}+\sum_{j\geq 1}\pi_{j}\sum_{i\geq 2}\bar{a}_{i}$ (39)
Substituting $\sum_{i\geq 1}\pi_{i}=1-\pi_{0}$ gives
$(1-\bar{a}_{1})(1-\pi_{0})=-\pi_{0}\beta+\pi_{0}\sum_{i\geq 1}\bar{a}_{i}+(1-\pi_{0})\sum_{i\geq 2}\bar{a}_{i}$ (40)
and moving terms, we have
$\alpha\pi_{0}$ $=1-\sum_{i\geq 1}\bar{a}_{i}$ (42)
$=1-2\beta-\sum_{i\geq 1}\sum_{j\geq i}\bar{\alpha}_{j}$ (43)
$=1-2\beta-\sum_{j\geq 1}\sum_{i=1}^{j}\bar{\alpha}_{j}$ (44)
$=1-2\beta-\sum_{j\geq 1}j\bar{\alpha}_{j}$ (45)
$=1-2\beta-\alpha\lambda\Delta$ (46)
gives the desired result. $\blacksquare$
We denote the upper bound on the lead as $\bar{L}$ and the PMF of $\bar{L}$ with $P^{\prime}_{1}$.